A solution containing 15.60g of ethanol in 108.0 g of H2O is prepared. Calculate the following Volumen .200L Explain the calculations


molar mass
volume of water
mass ethanol fraction
molarity of the solution
solution molality
molar fraction of water

molecular formula mass molar mass moles

solute
solvent
solutionProblema #4: (20) Se prepara una solución que contiene 15.60 g de etanol en 108.0 g de siguiente: etanol en 108.0 g de H20. C
Answer #1

Given Mass of Salute Cethanal) - 15.60g. Mas Solvent (Water) = 118.DOG Total mass of Salution - 15.60+108.00 = 123.60 g ( 123(6) Mass traction of = Ethanol 15.6 x 100% 15.6+ 108 = 12.62 % D Mass traction of = 108 x 200% Water 123.6 = 87.38% (8 MolarDensity of water is 1g/ml,(volume of solution considered as equal to mass of solution).